OpenStudy (anonymous):

Which ratio is equivalent to 6 : 30? A. 1 : 3 B. 1 : 5 C. 1 : 24 D. 1 : 6

OpenStudy (anonymous):

Both of the numbers in the ratio are divisible by 5. So, 6/6=? 30/6=?

OpenStudy (anonymous):

B.)

OpenStudy (anonymous):

Can you figure it out from there?

OpenStudy (anonymous):

1 : 5

OpenStudy (anonymous):

You guys should have tried to let them figure it out. :P

OpenStudy (anonymous):

gibson is the won who told her it

OpenStudy (anonymous):

/him

OpenStudy (anonymous):

Instead the first person to blurt out the answer got a medal, and the person who asked the question doesn't know how to answer it.

OpenStudy (anonymous):

her/him

OpenStudy (anonymous):

I dont care about the medal. I'm just trying to help the person that asked the question

OpenStudy (anonymous):

I'll explain it in case they still want to learn. Because both the numbers are divisible by 6, you would divide both by 6. 6/6=1 30/6=5 This is how you would get the ratio 1:5

OpenStudy (anonymous):

you cant just tell them the answer you have to tell them how to solve it

OpenStudy (anonymous):

so they can get the answer

OpenStudy (anonymous):

Just to be clear, I wasn't saying anything bad about any users. I was just putting my thoughts out as my own opinion and observation. I too don'tr care about the medals, it was just the principle.

OpenStudy (anonymous):

Ok. I will do that in the future. Thank you

OpenStudy (anonymous):

*don't

OpenStudy (anonymous):

I dont ether